OpenStudy (sugarmochi):

PLEASE PLEASE HELP. THIS IS URGENT Find the value of K if the lines kx-3y=4 and 4x-(k+1)y=7 are perpendicular

OpenStudy (drmoss):

The value of k is -3/7

OpenStudy (drmoss):

you good?

OpenStudy (sugarmochi):

how were you able to reach that answer..?

OpenStudy (drmoss):

y = (k/3)x - 4/3 -(k + 1)y = -4x + 7 y = [4/(k + 1)]*x - [7/(k + 1)] to be perpendicular the slopes must be negative reciprocals, then: 4/(k + 1) = -3/k 4k = -3k - 3 7k = -3 k = -3/7

OpenStudy (sugarmochi):

ohmygod thankyou so much! uhm uhm uhm is it possible for you to help me with another question?

OpenStudy (drmoss):

maybe

OpenStudy (drmoss):

I have to go soon because I'm on a tight schedule to finish my course.

OpenStudy (sugarmochi):

Find a,b if the line ax+by+3=0 passes through (0,1) and is parallel to 2x-3y+5=0 only help if you have time !

OpenStudy (drmoss):

i have to go

OpenStudy (drmoss):

be back ltr

OpenStudy (sugarmochi):

okay !

OpenStudy (pythagoras123):

\[ax+by+3=0\] Sub x = 0, y = 1 into the above equation: \[b+3=0\] \[b=-3\] For the line \[2x-3y+5=0\] can be re-written as \[y=\frac{ 2 }{ 3}x+\frac{ 5 }{ 3 }\] Sub b = -3 back into ax + by + 3 = 0 \[-3y=-ax-3\] \[y=\frac{ a }{ 3 }x+1\] Since the two lines are parallel, their gradients are equal. \[\frac{ a }{3 }=\frac{ 2 }{ 3 }\] By inspection of numerators, \[a=2\]